OpenStudy (anonymous):

Solve. e^2x+e^x-6=0 a. 0.69 b. 0.69, 1.10 c. 1.10, 0.14 d. 0.14

OpenStudy (anonymous):

SEX????

OpenStudy (anonymous):

@DarkgasherBeyblade What????

OpenStudy (anonymous):

Hey... do this... substitute e^x = t... You will get a quadratic equation in 't'... t^2 + t-6 = 0

OpenStudy (anonymous):

Ok, well I got the whole -3 and 2 answer but I dont know what to after that

OpenStudy (anonymous):

@saloniiigupta95

OpenStudy (anonymous):

All fine... now do this... e^x =-3 ==> x= ? e^x = 2 ==> x= ?

OpenStudy (anonymous):

Taking much time... hint: log for negative numbers is undefined...

OpenStudy (anonymous):

Lol so -3 is no solution? But how do I find the answer for 2?

OpenStudy (anonymous):

you will have to use log tables for that... You can definitely reject 'b' and 'c' between :-) Hope you know...  Use that to convert it into base 10 form... and log 2 (to the base 10) = 0.3011 approx.